# A new general quantum state verification protocol by the classical shadow method

### v3

@@ V.3.2 $n$-qubit GHZ state
For two different uniform sampling over all symplectic vectors $\mu$ with weight $t=(n-r)$ and all symmetric equivalence classes $[\mu]$, the second-largest eigenvalue of the level-$r$ strategy operator $\Omega$ always corresponds to the eigenvector $\mathcal{S}_{(1,\boldsymbol{0}^{n-1})}$ in the stabilizer basis of $|\mathrm{GHZ}_{n}\rangle$. For the two cases, the second-largest eigenvalues are $1-(2/3)^{n-r}$ and $(n-r)/(n-r+2)$, then the spectral gaps are
@@
The proof of this theorem is provided in the Appendix A. It’s astonishing to find that after considering the symmetries of $|\mathrm{GHZ}_{n}\rangle$, the spectral gap of the strategy operator for the uniform sampling is boosted from exponentially small to the inverse linear for $n$. The weight of Pauli measurements used in the level-$r$ protocol is $t=n-r$, then the spectral gap becomes $2/(t+2)$, which is independent of the qubit number $n$! We also see that increasing the level $r$ doesn’t improve the sample complexity.
@@ VI Conclusion
In this paper, we propose a new protocol—the DPSO protocol—which generalizes the ideas of the SOP introduced in [1] and [23]. The DPSO protocol inherits the key advantage of the SOP, namely the ability to verify almost all Haar-random pure states, while extending its applicability to a broader class of states in a more natural manner. Moreover, our protocol achieves lower sample complexity compared to the SOP. In particular, we demonstrate that the DPSO protocol exhibits a notably elegant structure when applied to stabilizer states, just as [8]. Besides, we also formulate the SOP in the form of hypothesis testing and discuss the relationships between the PLM framework and the SOP in details. The SOP deals with the problem of QSV in a new different view. In general, the sample complexity of the SOP is worse than that of the PLM framework with respect to the infidelity $\epsilon$ and the spectral gap $\nu(\Omega)$. But, it’s possible to verify many types of states uniformly instead of designing different strategy operators in the PLM framework for different types of states.
@@
Considering the current situation of the research of QSV, several challenges remain for the SOP or DPSO protocol to address in the future. First, it’s possible that this protocol can handle more special states, beyond stabilizer states to include non-stabilizer states. Second, it’s possible that the DPSO protocol with high level $r$ can also verify almost all Haar random pure states, as the level-$1$ SOP protocol [1]. Third, in the real-world experiments, the physical noise is unavoidable, then we need to consider the consequence of physical noise for the DPSO protocol. Hence, Addressing these challenges is an important direction for future research.
Acknowledgements
We thank Huangjun Zhu, Yunting Li, Zhihao Li, and Datong Chen for their helpful discussions, with special thanks to Yunting Li for providing the draft of her work with Huangjun Zhu.
Appendix A Proof of Theorem V.2
@@ Acknowledgements
We first consider the structure of the group $\Omega_{\mu}$ for different composite Pauli measurements with weight $t=n-r$. Suppose the Pauli measurements $\sigma_{l_{j_{1}}},\cdots,\sigma_{l_{j_{n-r}}}$ are applied to the set of qubits $J=\{j_{1},\cdots,j_{n-r}\}$, the left qubits are denoted by $K=\{k_{1},\cdots,k_{r}\}$.

### v4

@@ V.3.2 $n$-qubit GHZ state
For two different uniform sampling over all symplectic vectors $\mu$ with weight t=(nr)t=(n-r) and all symmetric equivalence classes $\mu]$, the second-largest eigenvalue of the level-$r$ strategy operator $\Omega$ always corresponds to the eigenvector $\mathcal{S}_{(1,\boldsymbol{0}^{n-1})}$ in the stabilizer basis of $|\mathrm{GHZ}_{n}\rangle$. For the two cases, the second-largest eigenvalues are $1-(2/3)^{n-r}$ and (nr)/(nr+2)(n-r)/(n-r+2), then the spectral gaps are
@@
The proof of this theorem is provided in the Appendix B. It’s astonishing to find that after considering the symmetries of $|\mathrm{GHZ}_{n}\rangle$, the spectral gap of the strategy operator for the uniform sampling is boosted from exponentially small to the inverse linear for $n$. The weight of Pauli measurements used in the level-$r$ protocol is t=nrt=n-r, then the spectral gap becomes $2/(t+2)$, which is independent of the qubit number nn! We also see that increasing the level $r$ doesn’t improve the sample complexity.
Compared with the known optimal protocols in the literature Pallister et al. (2018); Li et al. (2020); Dangniam et al. (2020), the sample complexity of our verification protocol is not optimal and grows linearly with the number of qubits $n$. However, this trade-off comes with the advantage that the protocol is easier to implement and can accommodate a broader class of quantum states simultaneously. There are also some papers addressing the verification of GHZ states from the perspectives of self-testing and cryptography Li et al. (2019); Hayashi and Koshiba (2022). These approaches are based on a fundamentally different assumption that the quantum devices or participants in the protocol may be untrusted, then the sample complexity of these protocols are much larger than our protocols, since we assumes that the measurement outcomes are fully reliable.
@@ VI Conclusion
In this paper, we propose a new protocol, called DPSO protocol, for verifying of a pure state under the assumption of the i.i.d., which generalizes the ideas of the SOP introduced in Huang et al. (2024a) and Li and Zhu (2025). The DPSO protocol inherits the key advantage of the SOP, namely the ability to verify almost all Haar-random pure states, while extending its applicability to a broader class of states in a more natural manner. Moreover, our protocol achieves lower sample complexity compared to the SOP. In particular, we demonstrate that the DPSO protocol exhibits a concise structure when applied to stabilizer states, just as Dangniam et al. (2020). Besides, we also formulate the SOP in the form of hypothesis testing and discuss the relationships between the PLM framework and the SOP in details. The SOP deals with the problem of QSV in a new different view. In general, the sample complexity of the DPSO (SOP) is worse than that of the PLM framework with respect to the infidelity $\epsilon$ and the spectral gap $\nu(\Omega)$. The DPSO protocol assumes the states produced by the quantum device is i.i.d., while the PLM framework can deal with more general scenarios where the states can be adversarial. However, the advantage of the DPSO protocol is that it’s possible to verify many types of states uniformly instead of designing different strategy operators in the PLM framework for different types of states.
@@
Considering the current situation of the research of QSV, several challenges remain for the SOP or DPSO protocol to address in the future. First, it’s possible that this protocol can handle more special states, beyond stabilizer states to include non-stabilizer states. Second, it’s possible that the DPSO protocol with high level $r$ can also verify almost all Haar random pure states, as the level-$1$ SOP protocol Huang et al. (2024a). Third, in the real-world experiments, the physical noise is unavoidable, then we need to consider the consequence of physical noise for the DPSO protocol. Fourth, it is also of interest to consider combining the SOP or DPSO protocols with self-testing techniques to address verification problems in scenarios similar to that discussed in Hayashi and Koshiba (2022). Hence, Addressing these challenges is an important direction for future research.
Acknowledgements: We thank Huangjun Zhu, Yunting Li, Zhihao Li, and Datong Chen for their helpful discussions, with special thanks to Yunting Li for providing the draft of her work with Huangjun Zhu.
Author Contributions: Xiaodi Li proposed the idea, performed the analytical and numerical computations, contributed to the interpretation of the results, and prepared the manuscript.
Conflicts of Interest: The authors declare no conflicts of interest.
Appendix A Classical description of target states
In the DPSO protocol, the post-measurement state is computed according to Eq.(28) of the manuscript. Two steps primarily determine the computational cost: the contraction of $|\boldsymbol{z}\rangle$ with $|\psi\rangle$, and the contraction of $|\widetilde{\phi}_{K,\boldsymbol{z}}\rangle$ with itself. Here, we take the matrix product state (MPS) as a representative example of tensor network states. Suppose the tensor network representation of $|\psi\rangle$ is
$|\psi\rangle=\sum_{b_{1},\dots,b_{n}}A^{b_{1}}_{\alpha_{1}}A^{b_{2}}_{\alpha_{1}\alpha_{2}}\cdots A_{\alpha_{N-1}}^{b_{N}}|b_{1}\rangle\otimes\cdots\otimes|b_{N}\rangle.$ (77)
Without loss of generality, we assume that Pauli measurements are applied to the first nrn-r qubits. The computation of the contraction $\langle\boldsymbol{z}|\psi\rangle$ involves contracting the vector corresponding to $|z_{i}\rangle$ with the $i$th tensor $A^{b_{i}}$ for $1\leq i\leq n-r$, followed by the contraction of the first nrn-r tensors. Therefore, the computational complexity is $\mathcal{O}((n-r)D^{2})$, where $D$ is the bond dimension. The second step of the contraction of $|\widetilde{\phi}_{K,\boldsymbol{z}}\rangle$ with itself is just the ordinary inner product of two MPS, whose complexity is $\mathcal{O}(rD^{3})$. Hence, we conclude that the computational complexity of obtaining the post-measurement state is $\mathcal{O}((n-r)D^{2}+rD^{3})$, implying that the computation is efficient for MPS.
Unfortunately, for other classes of states, computing the post-measurement state may not be efficient. In contrast, for projected entangled pair states (PEPS), the exact contraction of two PEPS is known to be computationally hard. Therefore, the computation of the post-measurement state in this case is also exponentially hard. For fermionic Gaussian states, the Jordan-Wigner transformation allows these states to be mapped to multi-qubit states. However, under this mapping, Pauli-$X$ and Pauli-$Y$ measurements become non-local operations and generally transform Gaussian states into non-Gaussian ones. Consequently, after applying such measurements, computing the post-measurement state of large-scale fermionic Gaussian systems may become computationally inefficient.
However, we can consider this problem from a different perspective. Our assumption of the DPSO protocol that there exists a classical description of the target state is equivalent to the assumption of the query model in the SOP. However, there exists more stronger assumptions, for example, the query model in Gupta et al. (2025). Here we can generalize the query model of the SOP to allow access to the amplitudes of $|\psi\rangle$ in the following expansion:
$|\psi\rangle=\sum_{s_{1},\cdots,s_{n}}c_{s_{1},\cdots,s_{n}}|s_{1}\rangle\otimes\cdots\otimes|s_{n}\rangle,$ (78)
where each $s_{i}$ corresponds to an eigenstate of the Pauli-$X$, $Y$, or $Z$ operator. Under this assumption, the first step—the contraction of $|\boldsymbol{z}\rangle$ with $|\psi\rangle$—requires $2^{r}$ queries to the model, resulting in a time complexity of $\mathcal{O}(2^{r})$. In the second step, computing the inner product of $|\widetilde{\phi}_{K,\boldsymbol{z}}\rangle$ with itself involves $2^{r}$ multiplications and summations of amplitudes, also yielding a time complexity of $\mathcal{O}(2^{r})$. Therefore, the overall time complexity for computing the post-measurement state in a multi-qubit system is $\mathcal{O}(2^{r})$.
Appendix B Proof of Theorem V.2
@@ Appendix A Classical description of target states
We first consider the structure of the group $\Omega_{\mu}$ for different composite Pauli measurements with weight t=nrt=n-r. Suppose the Pauli measurements $\sigma_{l_{j_{1}}},\cdots,\sigma_{l_{j_{n-r}}}$ are applied to the set of qubits $J=\{j_{1},\cdots,j_{n-r}\}$, the remaining qubits are denoted by $K=\{k_{1},\cdots,k_{r}\}$.